OpenStudy (anonymous):

How many solutions will this system of equations have? y = 3.5x - 3.5 y = -3.5x + 3.5

OpenStudy (anonymous):

only one solution ... it is just a zero.

OpenStudy (anonymous):

You are saying that y=0 x=1 And yes 1 solution

OpenStudy (anonymous):

Can you help me with another question?

OpenStudy (anonymous):

Sure. Post it here or close this one and open a new question

OpenStudy (anonymous):

A consistent system of equations is a system with __________. the same line parallel lines intersecting lines and lines that have the same slope intersecting lines and lines that have the same equation

OpenStudy (anonymous):

A consistent system of equations is system with one or more solutions. So which one do you think is the solution?

OpenStudy (anonymous):

@Andras is the last choice?

OpenStudy (anonymous):

Yes, well done

OpenStudy (anonymous):

thanks :) I have more questions though :( can you help me?

OpenStudy (anonymous):

@Andras

OpenStudy (anonymous):

Sure let me see

OpenStudy (anonymous):

Choose the correct graph of the given system of equations. y - 2x = -1 x + 3y = 4 graph of two lines, one with a positive slope and one with a negative slope, that intersect at the point (negative 1, 1) graph of two lines that intersect at the point (1, 1) graph of two parallel lines with positive slopes None of the above

OpenStudy (anonymous):

y=2x-1 positive slope y=-1/3x +4/3 negative slope Try to solve it now!

OpenStudy (anonymous):

@Andras so now solution?

OpenStudy (anonymous):

no solution* and is it the third one? @Andras

OpenStudy (anonymous):

It must have a solution

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

I know the answer but I want you to do some work too :)

OpenStudy (anonymous):

is it this?

OpenStudy (anonymous):

Yes master!

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

Ok one last question :-) @Andras If you want to help though

OpenStudy (anonymous):

Sure bring it on

OpenStudy (anonymous):

Ok I'm going to open another question

OpenStudy (anonymous):

@Andras